OpenStudy (anonymous):

Add these mixed numbers and reduce to the lowest terms. 3 3/12 + 2 4/6 A. 4 11/12 B. 5 11/12 C. 6 11/12

OpenStudy (liamschumm):

Hello there, Jaquan! Glad to help.

OpenStudy (liamschumm):

So, how would you evaluate 1 + 1/2?

OpenStudy (liamschumm):

Hint: 1 = 1/1.

OpenStudy (vianne):

Answer: = 39/12 + 16/6 = ((39 × 6) + (16 × 12)) / (12 × 6) = (234 + 192) / 72 = 426/72 = 71/12 = 5 11/12

OpenStudy (vianne):

@jaquan99 ?